OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, -2) and perpendicular to y =1/4 x + 2.

OpenStudy (anonymous):

\[Y-Y1=m(X-X1) ;then : y+1=-4(X+2) ;; then Y+1 =-4X-8\] the equation is "Y+4X=-9"